OpenStudy (anonymous):

Identify the Exponent Law shown in the example below. (a2)3 = a2·3 = a6 Power of a Product Power of a Power Product of Powers None of the Above

OpenStudy (jessiegonzales):

B. Power of a Power :)

OpenStudy (anonymous):

thanks

OpenStudy (jessiegonzales):

YW :) MEDAL?

OpenStudy (skullpatrol):

$$\Huge (a^2)^3 = a^{2·3} = a^6$$

OpenStudy (skullpatrol):

$$\Huge (a^m)^n = a^{m·n}$$

OpenStudy (skullpatrol):

$$\Huge (a^m)^n = a^m*a^m*...*a^m$$ $$\Huge=a^{m+m+...+m}$$ $$\Huge =a^{m*n}$$